A rectangular garden has dimensions of 18 feet by 13 feet. A gravel path of equal width is to be built around the garden. How wide can the path be if there is enough gravel for 516 square feet?
Answer by checkley71(8403)

(18+2x)(13+2x)-18*13=516
234+26x+36x+4x^2-234=516
4x^2+62x-516=0
(2x+43)(2x-12)=0
2x-12=0
2x=12
x=6 feet is the width of the path.
proof.
(18+12)(13+12)=750
18*13=234
750-234=516